OpenStudy (madmerc):

Which of the following represents ninth root of x to the seventh power in exponential form? x^9/7 x^7/9 9x^7 7x^9

OpenStudy (anonymous):

x^7/9

OpenStudy (madmerc):

thank you!!! can you help me with one more?

OpenStudy (anonymous):

sure :)

OpenStudy (madmerc):

Find the volume of a rectangular prism if the length is 4x, the width is 5x, and the height is x2 + x + 2. Use the formula V = l ⋅ w ⋅ h, where l is length, w is width, and h is height, to find the volume.

OpenStudy (anonymous):

V=4x*5x*(x2+x+2) V=20x2(x2+x+2) V=20x4+20x3+40x2

OpenStudy (madmerc):

ohhhhh okay!!! thank you so much! can you help me with just one more thing?

OpenStudy (anonymous):

ok

OpenStudy (madmerc):

Rationalize the denominator of \[\frac{ \sqrt{-49} }{ (7 - 2i) - (4 + 9i)}\]

OpenStudy (madmerc):

the choices are \[\frac{ -49 + 2i }{ -40 }\] \[\frac{ 49 + 21i }{ 58 }\] \[\frac{ 77 + 21i }{ -2 }\] \[\frac{ -77 + 21i }{ 130 }\]

OpenStudy (anonymous):

the answer is -77+21i/130 as \[\frac{ \sqrt{-49}}{ (7-2i)-(4+9i) }\] \[\frac{ \sqrt{-1}.\sqrt{49} }{ (7-2i)-(4+9i)}\] \[\sqrt{-1}\]=i \[\frac{ \sqrt{49}i }{ (7-2i)-(4+9i)}\] \[\frac{ 7i }{ 3-11i}\] \[\frac{ -77+21i }{ 130 }\]

OpenStudy (madmerc):

THANK YOU!!! I hadn't been able to figure that out at all!

OpenStudy (anonymous):

no problem :)